When a pendulum that is 0.5m long swings back and forth, its angular displacement, theta , in radians, from rest is given by theta= (1/4 sin( pi/2t), where t is the time , in seconds.
a) what is the period...I think it is 4??
b) At waht time is the pendulum first displaced by an angle of 0.1 radians? no idea

Answer by htmentor(1343)

The general form for the pendulum motion is theta(t) = A*sin(wt), where A is
the amplitude and w is the angular velocity.
So the amplitude is 0.25 and the angular velocity of the pendulum
is w = pi/2 rad/s.
a) The period, T is given by 2*pi/w = 4 s
b) 0.1 = 0.25*sin(pi/2*t) -> t = arcsin(0.1/0.25*2/pi) -> t = 0.262 s